OpenStudy (anonymous):

7(x + 6) + 10x Use the distributive property and combine like terms. 17x + 6 17 x + 42 23x 7x + 6 + 10x

OpenStudy (anonymous):

@movielover133

OpenStudy (anonymous):

17x+42

OpenStudy (anonymous):

you need to distribute the 7 to everything in the parenthesis

OpenStudy (anonymous):

9 + 2 + 7x^2 + 2xy^2 + 5x^2 Combine like terms to simplify the expression. Hint: ^ symbol stands for an exponent. 9 + 9x^2 + 2xy^2 20x^2 + 7xy^2 11 + 12x^2 + 2xy^2 11 + 7x^2 + 7xy^2

OpenStudy (anonymous):

11+12x^2+2xy^2

OpenStudy (anonymous):

1) 4^3= I think it is 64 2) 8^0 - 2^2 = I think it is 4 3) (-5) ^ 2 = @movielover133

OpenStudy (anonymous):

1-64

OpenStudy (anonymous):

anything to the power of 0 is gonna be 1 so 2)-3

OpenStudy (anonymous):

3)25 a negative number with an even exponent will be positive

OpenStudy (anonymous):

5^-2 Remember your rules about NEGATIVE EXPONENTS. 25 1/25 10 -10 8^2 • 8^? = 8^9 What is the missing exponent? @movielover133 thanks for all the help!!

OpenStudy (anonymous):

0.04

OpenStudy (anonymous):

and the missing exponent is 7

OpenStudy (anonymous):

when you have the same base(8) you add the exponents

OpenStudy (anonymous):

@movielover133 @eunalexis 5^-2 is 0.04???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

they just gave me this to pick out from 25 1/25 10 -10